If one wants to test whether there is a difference in mean scores before and after an exam, which will be an appropriate test to use?

a) one sample z test

b) one sample t test

c) paired t test
d) two sample independent t test

If X is a random variable following a Uniform(a,b) distribution, which option represents an approxi- mate shape of its density function?

a) straight vertical line between a and b

b) straight horizontal line between a and b

c) symmetric U shaped curve between a and b

3. LetX∼N(20,1). WhatisP(X>20)?

a) 0.25

b) 0.5

c) 0.75

d) 0.99

If the data is in the form of a frequency table, it is best described using a:

a) histogram

b) pie-chart

c) stem and leaf plot

d) boxplot
Answer #1

1. Answer: Option C paired t test

To compare difference between population means before and after an trial or experiment, we use paired t-test

2. Answer: Option C Symmetric U shaped curve between a and b

Since it is continuous uniform distribution with values lies between a and b, hence frequency curve will be U shaped with values symmetrically distributed between range a and b

3. Answer: Option B 0.5

X ~ N(20,1)

P(x>20) = P[(x - 20) / 1 > (20 - 20) / 1]=P(z > 0) = 1-P(z < 0) = 1 - 0.5 = 0.5

4. Answer: Option A Histogram

Histogram helps in identifying distribution of data represented in frequency table i.e whether data is unimodal (one mode value) or bimodal (exists two mode values) or symmetrical or skewed as well as if it is skewed then whether it is positively or negative skewed.